HETAV PATEL

Grade 11,

prove that a^2 + b^2 +c^2 = (a^2)(b^2) is not true for any a,b,c where a,b,c belongs to positive integers

prove that a^2 + b^2 +c^2 = (a^2)(b^2) is not true for any a,b,c where a,b,c belongs to positive integers

Grade:11

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

a^2 + b^2 +c^2 = (a^2)(b^2)

a2(1-b2)+b2+c2=0

This is not going to be zero provided a,b,c are all positive integers...

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sanjana rajendran
13 years ago

but how can we say that a^2(1-b^2)+b^2+c^2=0